# 3D Anderson localization of light in disordered systems of dielectric particles

### v1

@@ 4 Conclusions
[…]
Methods We performed numerical simulations using a self-developed light scattering code based on the open source DGTD Maxwell solver MIDG published by Tim Warburton [52]. The CST Microwave Studio software package based on the Time Domain Finite Integration Technique was used for a qualitative benchmarking the DGTD simulation results for simplified setups. In our code we implemented plane-wave and focused beam lights sources, boundary conditions and material description. Here, we use continuous and Gaussian pulses. A plane wave source is realized using the total field/scattered field technique. For simulations with a focused beam we implemented the plane wave superposition method [53]. An open computational domain is pretended with the PML absorbing boundary condition. For the time-resolved transmission measurements we illuminated target layers with a linearly polarized short plane wave pulse. The intensity of the transmitted signal is recorded, correspondingly, in a monitor plane behind the layer. The intensity curves in all plots are normalized by the peak intensity of the transmitted pulse. For generation of the particulate samples with different dimensions and controlled filling fractions we use the Bullet Physics engine [47]. This is an open-source C++ library that can be also used as an add-on in the Blender software. It allows simulation of dynamics and collision of multiple arbitrary 3D shapes in time domain. Thus, simulating natural powder mechanics samples with realistic topologies can be created. For sparse distributions of the pre-generated random irregular shapes we simulated their free fall on a substrate in a closed cylindrical volume. The sets of particles are created with a self-developed shape generator based on the Gaussian random field approach [46]. 3D models of the generated layers were then used for tetrahedral mesh generation. The spatial resolution in the regions with dielectric material with $n=3.0$ was $\sim 15$ tetrahedral cells per central wavelength of the incident pulse and the nodal expansion order in the DGTD numerical scheme was $N=3$. All the DGTD simulations were done on the HPC cluster Noctua 2 of Paderborn University using up to 30 cluster nodes per simulation.
Appendix A Supplementary material
A.1 Near field dynamics
Animated evolution of the near field in time helps to better represent the field dynamics in such a complex structure as a discrete random medium consisting of densely packed irregular particles. To support our analysis of the static field snapshots in Figure 1b we additionally simulate propagation of a short focused pulse in the same dense layer sample (Figure 1a) with the same model parameters. Animation A.1 represents an $XZ$ cross-section of the computational domain with the near-field intensity distribution and shows a cellular-automata-like behavior of the local field hotspots in the sub-wavelength voids between the closely neighboring particles.
Figure A.1: Animated propagation of a short focused pulse in a dense layer shown in Figure 1a.
Animation A.2 demonstrates dynamics of the near field that is trapped by the air-dielectric interface of the constituent particles after propagation of a short plane-wave pulse in a short period near $t=2t_{D}$. At high degree of disorder and in the absence of large voids the evanescent coupling mechanism slows down the energy decay from the system.
Figure A.2: Animated dynamics of the near field in a dense layer shown in Figure 1a excited by a short plane wave pulse.
A.2 Pulse transmission by layer samples with reduced dimensions
We take a cylindrically symmetric sample with $kR=26$, $kL=12$ and $\rho=0.44$ as a reference and consider reduced structures with thicknesses $kL=6$ and 3 and an elementary bi-layer structure. The refractive index is fixed at $n=3.0$. The results of the simulated transmission $T(t)$ for four layers are presented in Figure A.3a. A bi-layer shows purely exponential decay even though a WG mode is excited. The case of $kL=3$ is equivalent to a doubling the thickness of the bi-layer sample. This appears to be enough to make the $T(t)$ curve non-exponential. In all four cases a WG mode is excited allowing us reaching the $25-40t_{D}$ region even with such a small thickness. Further thickness increasing leads to stronger deviations of exponent and faster reduction of the diffusion coefficient $D(t)$ (Figure A.3b). However, only with $kL=12$ we obtain a qualitative change at longer times and reach the $t^{-1}$ fit.
Figure A.3: (a) Normalized transmission of a short pulse through layers with cylindric symmetry and different thickness $kL$ as a function of time $T(t)$. The diameter and volume fraction are $kR=26$ and $\rho=0.44$, correspondingly, and the refractive index is $n=3.0$. The energy decay from a simple double layer of particles is exponential while thicker layers demonstrate increasingly non-exponential transmission. (b) Diffusion coefficient $D(t)$ obtained by local exponential fitting of $T(t)$. $D(t)$ quickly evolves from a constant to a $t^{-1}$ fit (dashed line) with increasing thickness. (c) Vertical cross-sections of the samples with different thickness.
Figure A.4: Normalized transmission of a short pulse through a small layer of 2000 particles packed with volume fraction $\rho=0.5$ (inset) and with dimensions $kR=15$ and $kL=3$ as a function of time $T(t)$ at different refractive indices. Simulations are done with the CST Microwave Studio software package.
For the computations with the CST FIT solver we take a cylindric sample with thickness $kL=3$ (a four-particle thick layer) and reduced diameter $kR=15$. Due to parallelization restrictions of CST MWS we need to simplify the target system. With the particle size $kr=1$ and the volume fraction $\rho=0.5$ the number of particles becomes relatively small in this case, N=2000. With fixed sample parameters we vary the material refractive index $n$ from 2.0 to 3.0. For the CST computations we use a polyhedral description of the model sample and, then, a hexahedral mesh is generated by the software internally. The electromagnetic problem is solved in time domain simulating propagation of the short plane-wave pulse, similarly to our DGTD method setup.
The result in Figure A.4 is qualitatively similar to the $T(t)$ refractive index dependence in Figure 5a obtained for a much larger system with the DGTD method.
Summarizing all our results we note that increasing three parameters: layer thickness, volume fraction and refractive index act in very similar ways causing transition from an exponential dependence of $T(t)$ to a non-exponential behavior with qualitative change at longer times.

### v4

@@ III Analysis of the near field dynamics
First of all, we need to show that such a system reveals a non-exponential tail in $T(t)$ and a time-dependent diffusion coefficient $D(t)$ at longer times. Despite the small dimensions of the unit cell and small number of particles ($N=5000$ for $\rho=0.44$) we reach long enough observation times and see a qualitative transition in light transmission characteristics, very similar to that observed for larger finite cylindric samples. Figure S6 shows an emergence of a non-exponential tail in $T(t)$ at larger volume fraction ($\rho=0.44$) (a) and a time-dependent diffusion coefficient $D(t)$ (b). Transmission spectra for different volume fractions also show a transition from a number of overlapping peaks to isolated very sharp ones (Figure S6c). The horizontal axis in the spectral plots is scaled in the units of the particle size parameter, so that the pulse central wavelength corresponds to $X_{r}\approx 1.15$. Such a property of the dense system is sample-independent. We additionally simulated transmission for another realization of disorder for $\rho=0.44$ and it demonstrates qualitatively similar $T(t)$ dependence and transmission spectrum (Fig S7).
Figure S7: Normalized transmission $T(t)$ (a) and transmission spectra (b) for two realizations of disordered layers with volume fraction $\rho=0.44$.
Figure S8: Examples of the near-field intensity distributions $|E|^{2}$ at different time steps, shortly after $t=4t_{a}$ (a) and at $t>75t_{a}$ (b), when the transmission dependence $T(t)$ in Figure S6a becomes significantly non-exponential in the case of volume fraction $\rho=0.44$.
Figure S6d shows distributions of the near-field intensity in $XY$-plane for different volume fractions recorded at different times, when normalized transmission reaches $T(t)\approx 10^{-8}$. Here one can also notice a qualitative change. We see dipolar excitations of single particles and multiple wave scattering in a sparse structure, whereas, in a dense packing, the field energy is concentrated in small voids. Moreover, one can see clusters of hotspots and dark regions for the $\rho=0.44$ sample.
Now we can focus on the dense sample and track the changes in the near-field dynamics in time. Figure S8 represents field snapshots taken for the $\rho=0.44$ case at different time-steps before and after the $T(t)$ curve becomes non-exponential. The upper raw (Figure S8a) corresponds to the times shortly after $t=4t_{a}$ and shows randomly scattered hotspots, indicating diffusion. $XY$ and $XZ$ field maps in Figure S8b are taken at $t>75t_{a}$ when $T(t)$ is non-exponential and $D(t)$ is significantly time-dependent. At such long times field energy distribution turnes into a structured pattern of interacting clusters of hotspots. These clusters are, in fact, resonances induced by coherent propagation along highly disordered pathways. Video III shows their dynamics. The clusters (as well as dark regions) evolve in time but remain stable for tens of oscillation periods which is consistent with the sharp-peak spectrum. Single hotspots may have lifetimes of 10-20 periods. The lifetimes of the modes in the spectrum in Figure S6d ($\rho=0.44$) can be roughly estimated from the peak widths as long as $\sim 100$ periods which is a strong indication of localization.
To quantify this dynamics we calculated the inverse participation ratio (IPR) for such field maps for different volume fractions and different times using the $XY$-cross-section data. For a 2D scalar field $z(x,y)$ it is typically defined as
$\textrm{IPR}=\frac{\sum_{i}{z_{i}}^{2}}{{\left(\sum_{i}{z_{i}}\right)}^{2}}.$ (1)
$z$ is the near-field intensity in our case and IPR gives a measure of localization: lower IPR means energy spread out over many points and higher IPR implies localization at few points.
Figure S9: Time depence of the inverse participation ratio calculated from the intensity maps for different volume fractions. Linear regression indicates increasing energy concentration in loclaized modes with time. Dashed lines show a 95% confidence interval.
The calculated IPR as a function of time shows persistent growth for the volume fraction $\rho=0.44$ in Figure S9, that confirms our observation. Interestingly, this measure reflects more uniform spread of energy in sparse samples for the same measurement period. At $\rho=0.44$, field becomes increasingly concentrated in smaller spatial regions and the system evolves towards stronger spatial localization over time. The contrast between bright clusters of concentrated field hotspots and dark regions persistently grows. Therefore, such a process is not stationary and the observed resonances are not eigenmodes of the system but rather a time-dependent superposition of multiple localized modes. The excited modes can be initially isolated but their superposition and interference between them creates apparent spatial connections. Thus, we can conclude that we observe, for the first time, a real-time dynamics of Anderson localization formation in 3 dimensions.
We can give now a general interpretation of our results. Initial excitation by a short pulse creates superposition of many modes. Light explores many pathways, some of them begin to dominate due to interference while others decay. Obviously, extended modes decay faster then localized or quasi-localized ones. When the total energy of localized modes is significant enough, hotspot clusters become distinct and we obtain a slowdown in $T(t)$. After that, the system evolves towards pure Anderson modes with infinite lifetimes, which cannot be eventually reached due to the finiteness of the sample and its small thickness.
The localization length can be roughly estimated from the field images at different times as a few particle lengths or $\xi\approx 1-2\lambda$. This is also consistent with the layer thickness test in Figure S1. Individual modes can localize at the layer thickness $X_{L}=9.6$ ($L>\xi$) but the system is not large enough for efficient mode isolation ($L$ is not $\gg\xi$). Most of them extend across a significant fraction of the layer thickness which leads to spatial overlaps and interactions with the layer boundaries, increasing energy leakage. In fact, our simulation shows how Anderson localization manifests itself in a finite system. In order to directly observe a high-quality mode isolation we need a layer thickness of at least $L\sim 10\xi$ or $L\sim 10-20\lambda$ and with $\sim 10^{5}$ particles. This would require very large computing power, which is achievable, in principle, in large-scale massively parallel simulations. With our largest samples we get closer to this regime as it can be seen in Figure S1. Thus, it is not enough just to reach long observation times like we do this incorporating the WG-effect or periodic boundary conditions. Providing sufficiently large 3D volume for isolation of the disorder-induced localized modes is important.
We note, that the mechanism described above can be relevant also for low-absorbing materials as it involves mostly field-interface interaction and propagation in free space. The role of the real part of the complex refractive index becomes clear in this situation. It determines the balance between the energy transmitted by particles and groups of particles, enhancing a diffusive process, and the energy flowing through the random network of connected voids with long path lengths that enforces localization.